at a concession stand, five hot dogs and four hamburgers cost $13.00; four hot dogs and five hamburgers cost $14.00. Find the cost of one hot dog and the cost of one hamburger
Answer by lwsshak3(11628)

at a concession stand, five hot dogs and four hamburgers cost $13.00; four hot dogs and five hamburgers cost $14.00. Find the cost of one hot dog and the cost of one hamburger.
***
let x=cost of one hot dog
let y=cost of one hamburger
..
5x+4y=13
4x+5y=14
..
20x+16y=52
20x+25y=70
subtract
-9y=-18
y=2
4x=14-5y=14-10=4
4x=4
x=1
cost of one hot dog=$1.00
cost of one hamburger=$2.00